1. y'''+3y''+4y'-8y=0    ans, y= c1ex+e-2x(c2cos2x+c3sin2x)
  2.  y(4)-4y''=0                   ans, y=c1+c2x+c3e2x+c4e-2x
1
Expert's answer
2021-12-09T14:32:48-0500

1.

Corresponding (auxiliary) equation


"r^3+3r^2+4r-8=0"

"r^2(r-1)+4r(r-1)+8(r-1)=0"

"(r-1)(r^2+4r+8)=0"

"(r-1)((r+2)^2+4)=0"

"r_1=1, r_2=-2+2i, r_3=-2-2i"

The general solution of the given differential equation is


"y=c_1e^x+e^{-2x}(c_2\cos(2x)+c_3\sin(2x))"

"+c_3\sin(2x)+c_4x\sin(2x)"

2.

Corresponding (auxiliary) equation


"r^4-4r^2=0"

"r_1=r_2=0, r_3=2, r_4=-2"

The general solution of the given differential equation is


"y=c_1+c_2x+c_3e^{2x}+c_4e^{-2x}"